For the month of july you have planned sales of $59300 and decided a stock-to-sales ratio of 2.5 would be appropriate, while planned stock turn is 2.3. what should the BOM inventory be in order to support planned sales?
Answer #1

Let the BOM or initial stock level be I1 and the EOM or final stock level be I2

Average Inventory = AI = (I1+I2) / 2

Planned Sales = $ 59300, Stock to Sales Ratio = AI / Planned Sales = 2.5

AI = 59300 x 2.5 = $ 148250

I1 + I2 = 148250 x 2 = $ 296500 ------ (A)

Stock Turn = 2.3 = COGS / AI

COGS = 2.3 x 148250 = $ 340975

Now, I1 - COGS = I2

I1 - I2 = COGS = $ 340975 ---------- (B)

Adding Equation (A) and (B), we get:

2I1 = $ 637475

I1 = BOM = 637475 / 2 = $ 318737.5